The length of a rectangle is 23 centimeters less than six times its width. Its area is 35 square centimeters. Find the dimensions of the rectangle.
The width is ? cm.

Answer by jorel1380(3719)

w(6w-23)=35
6w2-23w-35=0
(6w+7)(w-5)=0
w=-7/6 or 5
Throwing out the negative answer, we get the width to be 5 cm., the length is 7 cm..